# System-Level Natural Language Feedback

### 25 Jan 2024 v2

@@ 1 Introduction
Figure 1: Our framework for incorporating NL feedback into system-level model design. Using a human-in-the-loop approach, criteria derived from NL feedback guide the creation of prompts for refining responses and metric design to evaluate the improvements. Notation: $q$: query, $r$: response, $f$: feedback, $r^{\prime}$: refinement, $m(\cdot)$: metric function.
@@ 5 Case Study 1: Query Generation
Table 2: Case study 1 (query generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for query refinement. NCR: non-copy rate, Spec.: specificity, Read.: readability, Con.: conciseness, Cov.: coverage. Sat.: satisfaction. The full criteria texts can be found in the Appendix.
@@ 5.1 Derive Criteria from Feedback
We collect all NL feedback from the FITS training split to understand human preferences and derive criteria. We first use SimCSE encoder444 We use the sup-simcse-roberta-large model. (Gao et al., 2021) to encode each feedback. Then, we use $k$-means clustering to group feedback related to query generation into five clusters. From inspecting these (see Appendix for detailed manual efforts), we summarize them into four groups (see Table 1) and derive that a successful search query should (i) rephrase the user’s question while keeping important keywords, (ii) be relevant and specific, (iii) use common words for better search coverage, (iv) be concise. The criteria text for crafting the prompt $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is in Table 2.
@@ 5.1.1 Criteria-guided Metric Design
  • Specificity measures whether the search query sufficiently captures the necessary information to retrieve relevant documents. We use GPT-3.5 as the evaluator (Fu et al., 2023). Details are in Appendix.
@@ 5.2.1 Refinement Generation
We use GPT-3.5 with criteria-based prompts to refine 1,000 randomly sampled unsatisfied queries (details in Appendix). To demonstrate the effectiveness of Figure 1-(A), we conduct ablation studies with different criteria for query refinement. Given our computational budget, for metrics relying on GPT-3.5, we sample 500 dialog contexts and compare the queries resulting from different criteria.
@@ 5.4 Evaluation using designed metrics
  • Results on Our Designed Metrics We also report results on our designed metrics for different query generators in Table 3. It is clear that training on satisfied data produces more specific and accurate queries, with further improvements when incorporating refinement data. The original BB2 query generator often generates overly concise queries, hindering the retrieval of the most relevant documents. In other words, although it generates queries that perform well in terms of readability or coverage, it is still an inadequate query generator, as evidenced by the poor satisfaction of the queries it generates. Later, when we refer to “our trained query generator”, we mean the one trained using both satisfied data and refinement data.
@@ 6 Case Study 2: Response Generation
Table 6: Case study 2 (response generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for response refinement. GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence. Sat.: satisfaction. The full criteria texts can be found in the Appendix.
@@ 6.1 Derive criteria from feedback
Following the approach in §5.1, we group all feedback related to response generation into ten clusters. Then, we summarize the following eight groups (see Table 5) of feedback types by merging some clusters. From Table 5, we derive that an improved response as indicated by users should (i) ground its answer on relevant search results, (ii) be concise and targeted, (iii) be confident in its answer. The exact criteria text for crafting the prompts $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 6.
@@ 6.1.1 Criteria-guided Metric Design
  • Factuality checks whether the information in the response is backed by search documents. We use GPT-3.5 with chain-of-thought to measure factuality (Luo et al., 2023). See Appendix for details.
  • Helpfulness measures whether the response directly answers the user’s question. We use GPT-3.5 to measure helpfulness. See Appendix for details.
  • Relevance measures whether the response remains on topic and offers pertinent information. We again use GPT-3.5, with further details in the Appendix.
@@ 6.2.1 Refinement Generation
We use GPT-3.5 with criteria-based prompts to refine 1,000 sampled unsatisfied responses (details in Appendix). As in §5.2.1, we conduct ablation studies to demonstrate the effectiveness of derived criteria. The results in Table 6 highlight: (i) Adding the groundedness criterion improves the groundedness metric. (ii) Adding the relevance criterion increases helpfulness and relevance. (iii) GPT-3.5 refinements are confident and rarely include phrases like “I’m not sure” or “I don’t know”. (iv) In terms of satisfaction, the best performance is achieved by the prompt with all criteria added. Therefore, when collecting training data, we use the three criteria-augmented prompt for response refinement.
@@ 6.4 Evaluation using designed metrics
  • SLT(QG)+BB2(RG) Original BB2 response model paired with our trained query generator.
@@
  • Results on Standard Metrics Standard metrics are shown in Table 8. Key takeaways include: (i) When using the BB2 response generator, our trained query generator improves the final response quality compared to the BB2 query generator. (ii) Training the response generator on satisfied data leads to further improvements when using our best query generator. (iii) However, training with additional refinement data does not surpass using satisfied data alone. The reason behind (iii) relates to FITS’s gold response collection. Often, the gold response is a user-guided, BB2-generated reply. This biases reference-based metrics towards original BB2 outputs. Moreover, low-quality references may underestimate model performance when using reference-based metrics (Zhang et al., 2023a) and we confirmed this with a human evaluation of response quality (see Appendix for details).
  • Results on Our Designed Metrics Table 7 shows the results when using our designed metrics. Notably, (i) when using BB2 response generator, our trained query generator improves the final response quality from all perspectives compared to the BB2 query generator. (ii) When equipped with our trained query generator, training the response generator on satisfied data leads to consistent improvements in helpfulness compared to the BB2 response generator, indicating the importance of domain-adapted training. (iii) Training the response generator on both satisfied and refinement data improves the final response quality from all perspectives compared to training on satisfied data only, highlighting refinement data’s utility in rectifying model errors. (iv) In terms of satisfaction, the best-performing system employs our query and response generators, both trained on satisfied and refinement data. Additionally, we gathered the first 200 unsatisfied responses into a sparse refinement training set, refined via instance-level feedback. A model trained on this set alongside satisfied data, fell short compared to our system-level trained response generator, as measured by our designed metrics (see Appendix for details).
@@ 7 Combining System-level Feedback and Instance-level Feedback
Previous studies (Scheurer et al., 2022b; Shi et al., 2022; Chen et al., 2023a) have shown the effectiveness of instance-level feedback in the refinement process. To take a step further, we explore the synergy of system-level and instance-level feedback on dialogue systems. Using response generation as a case study, we collect both human and GPT-3.5 feedback (prompt in Appendix) for the 1,000 unsatisfied responses from §6.2.1. We then design a refinement prompt integrating both system-level and instance-level feedback (see Appendix). We introduce three systems for comparison.
@@ 7.1 Results of Adding Instance-level Feedback
7.2 Compare Human and GPT-3.5 Feedback
@@
Figure 2: Compare human and GPT-3.5 feedback.
@@
In Figure 2, we show characteristics of human and GPT-3.5 feedback. Though GPT-3.5 feedback is lengthier and grammatically sound, it lacks the language diversity of human feedback. Upon manual examination, GPT-3.5 feedback is often general, whereas human feedback is direct and specific. See the Appendix for feedback examples.
@@ 7.3 Feedback Impact on Refinements
While GPT-3.5 feedback leads to more satisfactory refinements (see Figure 2), the performance of the resulting dialog system trained with these refinements falls short w.r.t. all our designed metrics compared to the system trained using human feedback-driven refinements as shown in Table 9. This disparity merits further scrutiny. Therefore, we also evaluate the refinement quality via designed metrics from §6.1.1, with results in Table 10. Refinements obtained using human feedback mainly stand out in groundedness and factuality. This aligns with the feedback clusters in Table 5 where over 40% of the feedback suggests the bot focus more on the search results; that is, focusing more on the search results will make the refinements more grounded, leading to a more grounded final system (see Table 9). Since language models are known to hallucinate regardless of their size (Ji et al., 2023; Li et al., 2023), grounding their generations to the documents is important to ensure factuality. Hence, groundedness of refinements plays an essential role in the performance of trained models.
@@ 8 Conclusion
In this paper, we present a framework that harnesses system-level NL feedback. By using a large set of instance-level feedback, we derive system-level feedback for refinement prompt engineering and metric design. We show the effectiveness of system-level feedback through two case studies: generating queries and formulating dialogue responses. We further combine system-level and instance-level feedback in the refinement data construction process, and observe that the resulting trained response generator makes considerable improvements versus using system-level feedback only. Finally, we explore the possibility of substituting instance-level human feedback with GPT-3.5 feedback. We find that human feedback stands out in capturing main issues, while GPT-3.5 feedback is lengthy and less focused.

### 03 Feb 2024 v3

@@ 1 Introduction
Figure 1: Our framework for incorporating NL feedback into system-level model design. Using a human-in-the-loop approach, criteria derived from NL feedback guide the creation of prompts for refining responses and metric design to evaluate the improvements. Notation: $q$: query, $r$: response, $f$: feedback, $r^{\prime}$: refinement, $m(\cdot)$: metric function. $S1\cdots S5$ represent different systems one can compare using this framework.
@@ 5 Case Study 1: Query Generation
Table 2: Case study 1 (query generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for query refinement. Metrics measured: NCR: non-copy rate, Spec.: specificity, Read.: readability, Con.: conciseness, Cov.: coverage. Sat.: satisfaction. The full criteria texts can be found in the Appendix A.2.
@@ 5.1 Derive Criteria from Feedback
We collect all NL feedback from the FITS training split to understand human preferences and derive criteria. We first use SimCSE encoder444 We use the sup-simcse-roberta-large model. (Gao et al., 2021) to encode each feedback. Then, we use $k$-means clustering to group feedback related to query generation into five clusters. From inspecting these (see Appendix A.1 for detailed manual efforts), we summarize them into four groups (see Table 1) and derive that a successful search query should (i) rephrase the user’s question while keeping important keywords, (ii) be relevant and specific, (iii) use common words for better search coverage, (iv) be concise. The criteria text for crafting the prompt $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is in Table 2.
@@ 5.1.1 Criteria-guided Metric Design
  • Specificity measures whether the search query sufficiently captures the necessary information to retrieve relevant documents. We use GPT-3.5 as the evaluator (Fu et al., 2023). Details are in the Appendix A.3.
@@ 5.2.1 Refinement Generation
We use GPT-3.5 with criteria-based prompts to refine 1,000 randomly sampled unsatisfied queries (details in Appendix A.2). To demonstrate the effectiveness of Figure 1-(A), we conduct ablation studies with different criteria for query refinement. Given our computational budget, for metrics relying on GPT-3.5, we sample 500 dialog contexts and compare the queries resulting from different criteria.
@@ 5.4 Evaluation using designed metrics
  • Results on Our Designed Metrics We also report results on our designed metrics for different query generators in Table 3. It is clear that training on satisfied data produces more specific and satisfactory queries, with further improvements when incorporating refinement data. The original BB2 query generator often generates overly concise queries, hindering the retrieval of the most relevant documents. In other words, although it generates queries that perform well in terms of readability or coverage, it is still an inadequate query generator, as evidenced by the poor satisfaction of the queries it generates. Later, when we refer to “our trained query generator”, we mean the one trained using both satisfied data and refinement data.
@@ 6 Case Study 2: Response Generation
Table 6: Case study 2 (response generation): refinement quality via designed metrics when using different criteria to prompt GPT-3.5 for response refinement. Metrics measured: GRD: groundedness, Fact.: factuality, Help.: helpfulness, Rel.: relevance, Conf.: confidence. Sat.: satisfaction. The full criteria texts can be found in the Appendix A.2.
@@ 6.1 Derive criteria from feedback
Following the approach in §5.1, we group all feedback related to response generation into ten clusters. Then, we summarize the following eight groups (see Table 5) of feedback types by merging some clusters. From Table 5, we derive that an improved response as indicated by users should (i) ground its answer on relevant search results, (ii) be concise and targeted, (iii) be confident in its answer. The criteria text for crafting the prompts $c$ for the text refiner $P_{\phi}(r^{\prime}|r,q,c)$ is given in Table 6.
@@ 6.1.1 Criteria-guided Metric Design
  • Factuality checks whether the information in the response is backed by search documents. We use GPT-3.5 with chain-of-thought to measure factuality (Luo et al., 2023). See Appendix A.3 for details.
  • Helpfulness measures whether the response directly answers the user’s question. We use GPT-3.5 to measure helpfulness. See Appendix A.3 for details.
  • Relevance measures whether the response remains on topic and offers pertinent information. We again use GPT-3.5, with further details in the Appendix A.3.
@@ 6.2.1 Refinement Generation
We use GPT-3.5 with criteria-based prompts to refine 1,000 sampled unsatisfied responses (details in Appendix A.2). As in §5.2.1, we conduct ablation studies to demonstrate the effectiveness of derived criteria. The results in Table 6 highlight: (i) Adding the groundedness criterion improves the groundedness metric. (ii) Adding the relevance criterion increases helpfulness and relevance. (iii) GPT-3.5 refinements are confident and rarely include phrases like “I’m not sure” or “I don’t know”. (iv) In terms of satisfaction, the best performance is achieved by the prompt with all criteria added. Therefore, when collecting training data, we use the three criteria-augmented prompt for response refinement.
@@ 6.4 Evaluation using designed metrics
  • SLT(QG)+BB2(RG) Original BB2 response model paired with our system level trained query generator.
@@
  • Results on Standard Metrics Standard metrics are shown in Table 8. Key takeaways include: (i) When using the BB2 response generator, our trained query generator improves the final response quality compared to the BB2 query generator. (ii) Training the response generator on satisfied data leads to further improvements when using our best query generator. (iii) However, training with additional refinement data does not surpass using satisfied data alone. The reason behind (iii) relates to FITS’s gold response collection. Often, the gold response is a user-guided, BB2-generated reply. This biases reference-based metrics towards the original BB2 outputs. Moreover, low-quality references may underestimate model performance when using reference-based metrics (Zhang et al., 2023a) and we confirmed this with a human evaluation of response quality (see Appendix A.4 for details).
  • Results on Our Designed Metrics Table 7 shows the results when using our designed metrics. Notably, (i) when using the BB2 response generator, our trained query generator improves the final response quality from all perspectives compared to the BB2 query generator. (ii) When equipped with our trained query generator, training the response generator on satisfied data leads to consistent improvements in helpfulness compared to the BB2 response generator, indicating the importance of domain-adapted training. (iii) Training the response generator on both satisfied and refinement data improves the final response quality from all perspectives compared to training on satisfied data only, highlighting refinement data’s utility in rectifying model errors. (iv) In terms of satisfaction, the best-performing system employs our query and response generators, both trained on satisfied and refinement data. Additionally, as a further baseline, we gathered the first 200 unsatisfied responses into a sparse refinement training set, refined via instance-level feedback. A model trained on this set alongside satisfied data, fell short compared to our system-level trained response generator, as measured by our designed metrics, see Appendix A.5 for details.
@@ 7 Combining System-level Feedback and Instance-level Feedback
Previous studies (Scheurer et al., 2022b; Shi et al., 2022; Chen et al., 2023a) have shown the effectiveness of instance-level feedback in the refinement process. To take a step further, we explore the synergy of system-level and instance-level feedback on dialogue systems. Using response generation as a case study, we collect both human and GPT-3.5 feedback (prompt in Appendix A.6) for the 1,000 unsatisfied responses from §6.2.1. We then design a refinement prompt integrating both system-level and instance-level feedback, i.e. both the desired criteria and the specific example-based feedback (see Appendix A.2). We introduce three systems for comparison.
@@ 7.1 Results of Adding Instance-level Feedback
7.2 Human vs. GPT-3.5 Feedback Metrics
@@
Figure 2: Comparison of human and GPT-3.5 feedback.
@@
In Figure 2, we show characteristics of human and GPT-3.5 feedback. Though GPT-3.5 feedback is lengthier and grammatically sound, it lacks the language diversity of human feedback. Upon manual examination, GPT-3.5 feedback is often general, whereas human feedback is direct and specific. See the Appendix A.7 for feedback examples.
@@ 7.3 Feedback Impact on Refinements
While GPT-3.5 feedback leads to a higher refinement success rate (see Figure 2), the performance of the resulting dialog system trained with these refinements falls short w.r.t. all our designed metrics compared to the system trained using human feedback-driven refinements as shown in Table 9. Therefore, to understand this further we also evaluate the refinement quality via designed metrics from §6.1.1, with results in Table 10. Refinements obtained using human feedback mainly stand out in groundedness and factuality. This aligns with the feedback clusters in Table 5 where over 40% of the feedback suggests the bot focus more on the search results; that is, focusing more on the search results will make the refinements more grounded, leading to a more grounded final system (see Table 9). Since language models are known to hallucinate regardless of their size (Ji et al., 2023; Li et al., 2023), grounding their generations to the documents is important to ensure factuality. Hence, groundedness of refinements plays an essential role in the performance of trained models.
@@ 8 Conclusion
In this paper, we present a framework that harnesses system-level NL feedback. By using a set of instance-level feedback, we derive system-level feedback for refinement prompt engineering and metric design. We show the effectiveness of system-level feedback through two case studies: generating queries and formulating dialogue responses. We further combine system-level and instance-level feedback in the refinement data construction process, and observe that the resulting trained response generator makes considerable improvements versus either alone. Finally, we explore the possibility of substituting instance-level human feedback with GPT-3.5 feedback. We find that human feedback stands out in capturing main issues, while GPT-3.5 feedback is lengthy and less focused.
@@ 9 Limitations
10 Acknowledgement
The work was done as part of the Meta–NYU mentorship program and partly supported by the National Science Foundation (under NSF Award 1922658). Kyunghyun Cho is supported by the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI).